# History of $N(1680)$

### v2

@@ Abstract
This paper describes my personal appreciation for some of the great research achievements of Mitya Diakonov, Vitya Petrov, and Maxim Polyakov and how my own research career has followed the paths they opened. Among the topics where they have been the most influential have been the pursuit and study of the exotic pentaquark. The search for exotics may require a complementary approach, such as experimental and theoretical activity. Here, I would like to focus on a story of $N(1680)$ in which I was involved.
@@ 1 Prehistory
I don’t remember Maxim well from the Leningrad Nuclear Physics Institute (LNPI and PNPI now) Winter School on Nuclear and Particle Physics times in the 1970s when he was a student. But somewhere in 1997 or so (the time of the famous $\overline{\textbf{10}}$ paper [1] came up), Dick Arndt forwarded me a letter from Maxim asking what the SAID kaon-nucleon Partial-Wave Analysis (PWA) [2] gives for the exotic $S=+1$ baryon. Dick was very surprised, since the analysis gave nothing which may fit such a narrow exotic state (all four resonances in standard PWA are too heavy $\mathrm{ReW}>1790~{}\mathrm{MeV}$ and too broad $\mathrm{|ImW|}>95~{}\mathrm{MeV}$, then so do not fit the predicted $\Theta^{+}$ in Ref. [1]). From that moment on, my friendship and long-term cooperation with Maxim began. To solve a PWA problem, we came up with a modified PWA to look for the narrow resonance [3].
@@
I met Mitya for the first time at one of the LNPI Winter Schools in the mid 1970s. It is known that the mass difference in multiplets is determined by the phenomenological Gell-Mann-Okubo mass formula, which depends on the nucleon sigma term. Our SAID sigma term [4] was selected for this purpose [1]. In 2005 or so, Mitya gave a talk at a JLab seminar and Dick Arndt, who was sitting next to me, asked where this boy got such English. My answer was simple - his mother (Nina Yakovlevna) translated a poem by John Keats from English into Russian. Dick was satisfied with my answer.
@@
I also met Vitya for the first time at one of the LNPI Winter Schools in the 1970s, but there is a background. In 1968, I measured the efficiency of neutron detection with large scintillation counters, using a beam of neutrons from the Neutron Generator (that was my MS project). At the same time, a physicist and an engineer studied the reaction of transistors to neutrons using the same source of fast neutrons. The physicist’s name was Tamara Markovna. One day, a strange man appeared looking for his wife. Somebody told me who he was, Yuri Viktorovich Petrov, a well-known theoretician. That’s how I found out Tamara Markovna’s last name (they were Vitya’s parents), and Vitya was maybe 13 years old at the time.
@@
Obviously, the Instanton Vacuum model is a very powerful tool [5]. Before the $\overline{\textbf{10}}$ tasks came up, our friends focused on the nucleon-nucleon interaction and, in particular, on the dibaryon (see, for instance, Ref. [6]). In 1997, my postdoc, Serezha Pavlenko, started to work with Vitya and Mitya. He was a brilliant young experimentalist and phenomenologist. In a short time, Serezha did many calculations and Mitya and Vitya had a high respect for Serezha. Unfortunately, a stroke killed him. In 1998, he was 24 years old.
@@ 2 Introduction
QCD gives rise to the hadron spectrum [7] and many $q\bar{q}$ and $qqq$ have been observed [8]. However, $q\bar{q}q\bar{q}$ and $qqqq\bar{q}$ and other many-quark states are not forbidden but have not yet been observed. Recently, the LHCb Collaboration claimed evidence for four hidden-charm $qqqc\bar{c}$ states near open-charm decay thresholds for $\Sigma_{c}^{+}\bar{D}^{0}$ and $\Sigma_{c}^{+}\bar{D}^{\ast 0}$ in $\Lambda_{b}\to J/\psi pK^{-}$ decay [9]. Nevertheless, although there is no doubt about the LHCb observations, it is not clear whether they are compact multi-quark or hadonic molecular states [10].
In the light quark sector, Diakonov, Petrov, and Polyakov proposed a clearly exotic and narrow $\Theta^{+}(uudd\bar{s})$ state [11], yet to be unequivocally observed and identified. The original name of the pentaquark lying at the apex of $\overline{\textbf{10}}$ was $Z^{+}$, then following Mitya’s suggestion, we called it $\Theta^{+}$ now (see Appendix). Using the Chiral Quark Soliton model [11], this particle along with other members of $\overline{\textbf{{10}}}$ has been proposed with a mass of $M_{\Theta^{+}}=1.53~{}\mathrm{GeV}$ and a width less than $15~{}\mathrm{MeV}$ [1] (Fig. 1). Due to relatively low mass and simple decay channels to $K^{+}n$ or $K^{0}p$, it has attracted attention of many experiments at different facilities worldwide (the Citation Index (CI) on this renowned paper [1] is about 1000).
Figure 1: The suggested anti-decuplet of baryons [1]. The three corners of the triangle (in red) are exotic, which means that their quantum numbers require more than three quarks.
@@
The standard PWA (by construction) tends to miss a narrow resonance with $\Gamma_{R}<20~{}\mathrm{MeV}$ or so [3]. Modified PWA assumes the existence of a narrow resonance by inserting a Breit-Wigner (BW) structure to one of the partial amplitude and refitting the whole database. Such an analysis of the effect of the scanning mass $M_{R}$, full $\Gamma_{R}$ and partial width of the narrow resonance for each partial wave allows us to get a picture. The worse description of the full database that includes resonance with the corresponding $M_{R}$ and $\Gamma_{R}$ (in terms of $\Delta\chi^{2}$) does not support our hypothesis. The better description may result that the resonance may exist, or the effect can be due to various corrections (e.g. thresholds), or both possibilities can contribute. Some additional checks are necessary: true resonance should provide the effect only in a particular partial wave, while a non-resonance source may show similar effects in various partial waves.
Using this kind of modified PWA, Arndt and co-workers performed a reanalysis of the existing $KN$ database [2] in order to refits $K^{+}N$ observables and consider the effect of a narrow state [3].
It was found that the existence of a $\Theta^{+}$ in the $P_{01}$ ($J^{P}=1/2^{+}$) state, with a mass of $\sim 1545~{}\mathrm{MeV}$ and a width of $\Gamma(\Theta^{+})\leq 0.5~{}\mathrm{MeV}$ or so is possible (Fig. 2).
Figure 2: Change of overall $\chi^{2}$ due to insertion of a resonance into $P_{01}$ for $M_{R}=1520–1560~{}\mathrm{MeV}$ with $\Gamma_{R}=0.3$ and $0.7~{}\mathrm{MeV}$, using KN plane wave approximation [2]. Energies, where measurements exist, are labeled by red vertical bars (for references, see Ref. [2]). The $1535~{}\mathrm{MeV}$ values for $\Delta\chi^{2}$ are off-scale [3]. The lines are given to guide the eye.
@@ 3 My Story of $N(1680)$
The mass difference in baryon multiplet is determined by the phenomenological Gell-Mann-Okubo mass formula (mixing is able to shift some masses). Originally, Mitya, Vitya and Maxim claimed $N(1710)$ as a non-strange member of $\overline{\textbf{10}}$, which was a suitable candidate for $N^{\ast}$ [14]. The state $N(1710)$, though listed in the PDG2002 baryon summary table [15] as a three-star resonance, was not seen in the PWA of the $\pi N$ elastic scattering data by the SAID group [16].
@@
Figure 3: Argand plots for the $P$ partial-wave amplitudes for the $\pi N$ case from threshold ($1080~{}\mathrm{MeV}$) to $W=2500~{}\mathrm{MeV}$ [17]. The crosses indicate $50~{}\mathrm{MeV}$ steps in $W$. Solid circles correspond to the BW $W_{R}$.
The $\pi$N PWA by the GW SAID group has shown that above the Roper resonance, $W>1500~{}\mathrm{MeV}$, $\sigma^{tot}\cong 2\sigma^{el}\cong 2\sigma^{inel}$ ($\sigma^{tot}=\sigma^{el}+\sigma^{inel}$). It means that is nearly pure diffraction, inelasticity $\eta\to\infty$ for $P_{11}$ and $S\cong 0$ with $A\cong i/2$. The result is that phase $\delta$ is poorly defined and the $P_{11}$ amplitude is spinning around $ReA=0$ [17] (Fig. 3).
Of course, the non-observation of a broad $N(1710)$ state in $\pi N$ elastic analyses (no pole position, no BW, no speed plot) could be due to a very small $\pi N$ branching ratio. Standard procedure used in PWA may also miss narrow resonances with $\Gamma<20~{}\mathrm{MeV}$ by construction [16]. Therefore, the true unitary partner of $\Theta^{+}$ (if it is different from $N(1710)$ and sufficiently narrow) could have eluded detection.
Arndt and co-workers were considering the $\pi N$ partial wave $P_{11}$, as this is amplitude, which is associated with resonances having $J^{P}=1/2^{+}$ [18]. The character of $\chi^{2}$ changes, $\Delta\chi^{2}$, after inserting a narrow resonance into the partial amplitude with a range of masses, widths, and branching fractions is illustrated in Fig. 4. Negative values of $\Delta\chi^{2}$ emerge more readily near $M_{R}=1680~{}\mathrm{MeV}$ and $1730~{}\mathrm{MeV}$. We see that $\Delta\chi^{2}$ becomes negative only for $\Gamma_{el}=(\Gamma_{el}/\Gamma_{tot})\cdot\Gamma_{tot}$ within the bounds $\Gamma_{el}\leq 0.5~{}(0.3)~{}\mathrm{MeV}$ for $M_{R}=1680~{}(1730)~{}\mathrm{MeV}$ (Fig. 4). The available data cannot reliably discriminate values of $\Gamma_{el}$ below these bounds.
@@
(Fig. 5 (left)) [19]. BTW, in the case of a “neutron” target, Final State Interaction (FSI) is a problem. The effect is absent in the case of $\gamma p\to\eta p$ (Fig. 5 (right)) and agrees with the prediction of the Chiral Quark Solution approach [20]. In exact SU(3)F, the transition magnetic moment $\mu(p^{\ast}\to p)$ should vanish, since the U-spins are 3/2 for $p^{\ast}$, 1/2 for $p$, and 0 for the photon. With violation of SU(3)F, this transition moment does not vanish but is still much smaller than $\mu(n^{\ast}\to n)$ [21].
@@
The GW SAID group performed a coupled channel analysis of the $\pi N$ system including $\pi^{\pm}p$ elastic scattering and $\pi^{-}p\to\pi^{0}n$ with $\pi^{-}p\to\eta n$ [16]. The attractive factor for the reaction $\pi^{-}p\to\eta n$ is that it plays a role in isospin filtering. Unfortunately, the $\pi^{-}p\to\eta n$ data above $800~{}\mathrm{MeV}$ ($W=1630~{}\mathrm{MeV}$) are not reliable for PWA (see, for instance, Fig. 6 (left)). Most of NIMROD data do not satisfy requirements (systematics is 10% or more, momentum uncertainties are up to $100~{}\mathrm{MeV/c}$, and so on). That is the main reasons why the SAID coupled channel analysis had no chance to look for $N(1680)$ using reaction data $\pi^{-}p\to\eta n$ (Fig. 6 (right)).
Figure 6: Left: Excitation function for differential cross sections for the reaction $\pi^{-}p\to\eta n$. Data came from BNL, RHEL, and Saclay. Solid blue line from SAID PWA (WI08 solution) [24]. Right: Total cross sections for the reaction $\pi^{-}p\to\eta n$. Red solid line from SAID PWA [24]. All data are available at Ref. [23].
EPECUR Collaboration at ITEP performed high quality measurements of the differential cross sections for the elastic $\pi^{\pm}p$ over an energy range of $p=800-1300~{}\mathrm{MeV/c}$ ($W=1.55-1.83~{}\mathrm{GeV}$) and for center-of-mass angles from 40 to $120~{}\mathrm{deg}$ [25]. In total, approximately 10,000 new data points have been obtained. These data have been produced with a momentum resolution of $\sim 1~{}\mathrm{MeV/c}$ and with $\sim 1$% statistical uncertainties.
Figure 7: Comparison of the measured $\pi$N data [23] and calculated $\pi^{-}p\to\eta n$ total cross sections. The red solid curve presents the calculations from the work [26]. The black dashed curve indicates the contribution of the $S$ wave.

### v3

@@ Abstract
This paper describes my personal appreciation for some of the great research achievements of Mitya Diakonov, Vitya Petrov, and Maxim Polyakov and how my own research career has followed the paths they opened. Among the topics where they have been the most influential have been the pursuit and study of the exotic pentaquark. The search for exotics may require a complementary approach, such as experimental and theoretical activity. Here, I would like to focus on a story of $N(1680)$, a non-strange unitary partner of $\Theta^{+}$, in which I was involved.
@@ 1 Prehistory
I am an experimentalist and served in the Leningrad Nuclear Physics Institute (LNPI or PNPI now) High Energy Physics Division (HEPD) for 29 years. In particular, the reaction $\pi p\to p\pi\pi$ was an important part of the HEPD research. From my friends Senya Sherman (experimentalist) and Tolya Bolokhov (theorist), I had the opportunity to learn about an extraordinary student, Maxim, who was heavily involved in this activity, and he had several joint papers with seniors back to the end of 1980s.
@@
Somewhere in 1997 or so (the time of the famous $\overline{\textbf{10}}$ paper [1]), Dick Arndt forwarded me a letter from Maxim asking what the Virginia Tech Scattering Analysis Interactive Dial-in (SAID) Kaon-nucleon Partial-Wave Analysis (PWA) [2] gives for the exotic $S=+1$ baryon. Dick was very surprised, since the analysis gave nothing which may fit such a narrow exotic state. All four resonances in standard PWA are too heavy (mass $>1790~{}\mathrm{MeV}$) and too broad (width $>95~{}\mathrm{MeV}$). So they do not fit the predicted $\Theta^{+}$ in [1]. This fact was a motivation for us to solve a standard PWA problem, and we do a modified PWA to look for narrow resonances [3].
It was a good surprise for me to see that the famous $\overline{\textbf{10}}$ paper [1] used our SAID sigma-term result [4] for the mass difference in anti-decuplet using the Gell-Mann-Okubo phenomenological mass formula.
The experimental results published in the top peer-review journals had problems for experimentalists and the main reason was English. I was a junior researcher at PNPI and was willing to write papers on $\pi^{\pm}p$ elastic scattering and proton-induced pion production to report the results of our group measurements at Gatchina. I translated the Russian versions into English and asked Mitya to polish them. Sure, Mitya ignored my silly staff, and one eve per paper was enough for him to make drafts which Nucl. Phis. A accepted in 1981 and then J. Phys. G accepted in 1988. It will be good to acknowledge Mitya for his help in translation, but our experimental physics was far away from his mainstream, and for that reason Mitya’s name is still hidden in both papers.
@@
In 2005 or so, Mitya gave a talk at a JLab seminar about pentaquark, and Dick Arndt, who was sitting next to me, asked where this guy got such English. My answer was simple - his mother (Nina Yakovlevna) was a professor of English literature in St. Petersburg State University and translated a poem by John Keats from English into Russian. Dick was satisfied with my answer.
@@
I met Vitya for the first time at one of the LNPI Winter Schools at the end of the 1970s, but there is a background. In 1968, I measured the efficiency of neutron detection with large scintillation counters, using a beam of neutrons from the Neutron Generator (that was my Master project). At the same time, a physicist and an engineer studied the reaction of transistors to neutrons using the same source of fast neutrons. The physicist’s name was Tamara Markovna. One day, a strange man appeared looking for his wife. Somebody told me who he was, Yurii Viktorovich Petrov, a well-known theoretician. That’s how I found Tamara Markovna’s last name. They were Vitya’s parents and Vitya was probably 13 years old at the time.
@@
Obviously, the Instanton vacuum model developed by Mitya and Vitya is a very powerful model [5]. Before the $\overline{\textbf{10}}$ tasks came up, our friends focused on the nucleon-nucleon interaction and, in particular, on the dibaryon (see, for instance, Ref. [6]). In 1996 or 1997, I asked my postdoc, Serezha Pavlenko, to work with Mitya and Vitya because that was a time when many groups were looking for dibaryon. He was a brilliant young experimentalist and phenomenologist. In a short time, Serezha did many calculations and Mitya and Vitya were impressed with the Serezha results. Unfortunately, he passed away shortly after the stroke. In 1998, Serezha was 24 years old.
@@ 2 Introduction
QCD gives rise to the hadron spectrum [7] and many $q\bar{q}$ and $qqq$ have been observed [8]. However, $q\bar{q}q\bar{q}$ and $qqqq\bar{q}$ and other many-quark states are not forbidden but have not yet been observed. Recently, the LHCb Collaboration claimed evidence for four hidden-charm $qqqc\bar{c}$ states near open-charm decay thresholds for $\Sigma_{c}^{+}\bar{D}^{0}$ and $\Sigma_{c}^{+}\bar{D}^{\ast 0}$ in $\Lambda_{b}^{0}\to P^{+}_{c\bar{c}}K^{-}\to J/\psi pK^{-}$ decay [9]. However, although there is no doubt about the LHCb observations, it is not clear whether they are compact multiquark or hadonic molecular states [10].
In the light-quark sector, Mitya, Vitya, and Maxim proposed a clearly exotic and narrow $\Theta^{+}(uudd\bar{s})$ state [11], yet to be unequivocally observed and identified. The original name of the pentaquark lying at the apex of $\overline{\textbf{10}}$ was $Z^{+}$, then following Mitya’s suggestion, we now call it $\Theta^{+}$ (see the Appendix). Using the Chiral Quark Soliton model [11], this particle along with other members of $\overline{\textbf{{10}}}$ has been proposed with a mass of $M_{\Theta^{+}}=1.53~{}\mathrm{GeV}$ and a width less than $15~{}\mathrm{MeV}$ [1] (Fig. 1). Due to the relatively low mass and simple decay channels to $K^{+}n$ or $K^{0}p$, it has attracted the attention of many experiments at different facilities around the world (this renowned paper [1] has about 1000 citations collected by the Citation Index (CI) at iNSPIRE HEP).
Figure 1: The suggested anti-decuplet of baryons [1]. A weight diagram for anti-decuplet formed by using two flavor conjugate diquarks $(\bar{q})_{qq}$ and one anti-quark $\bar{q}$ [12]. Here $q\bar{q}$ denotes $u\bar{u}$, $d\bar{d}$, and $s\bar{s}$. The three corners of the triangle (in red) are exotic, which means that their quantum numbers require more than three quarks.
@@
The standard PWA (by construction) tends to miss a narrow resonance with $\Gamma_{R}<20~{}\mathrm{MeV}$ or so [3]. The modified PWA [3] assumes the existence of a narrow resonance by inserting a Breit-Wigner (BW) structure into one of the partial amplitudes and refitting the whole database. Such an analysis of the effect of the scanning mass $M_{R}$, full $\Gamma_{R}$ and partial width of the narrow resonance for each partial wave allows us to get a picture. The worse description of the full database that includes resonance with the corresponding $M_{R}$ and $\Gamma_{R}$ (in terms of $\Delta\chi^{2}$) does not support our hypothesis. The better description may result in the resonance that may exist, or the effect can be due to various corrections (e.g., thresholds), or both possibilities can contribute. Some additional checks are necessary: true resonance should provide the effect only in a particular partial wave, while a non-resonance source may show similar effects in various partial waves.
Using this kind of modified PWA, Arndt and co-workers performed a reanalysis of the existing $KN$ database [2] to refit $K^{+}N$ observables and consider the effect of a narrow state [3]. The existence of a $\Theta^{+}$ in the $P_{01}$ ($J^{P}=1/2^{+}$) state, with a mass of $\sim 1545~{}\mathrm{MeV}$ and a width of $\Gamma(\Theta^{+})\leq 0.5~{}\mathrm{MeV}$ or so, was found to be possible (Fig. 2).
Figure 2: Change of overall $\chi^{2}$ due to insertion of a resonance into $P_{01}$ for $M_{R}=1520–1560~{}\mathrm{MeV}$ with $\Gamma_{R}=0.3$ and $0.7~{}\mathrm{MeV}$, using KN plane wave approximation [2]. Energies, where measurements exist, are labeled by red vertical bars (for references, see [2]). The $1535~{}\mathrm{MeV}$ values for $\Delta\chi^{2}$ are off-scale [3]. Lines are given to guide the eye.
@@ 3 My Story of $N(1680)$
The mass difference in multiplet is determined by the phenomenological Gell-Mann-Okubo mass formula (mixing is able to shift some masses). Originally, Mitya, Vitya and Maxim claimed $N(1710)$ as a non-strange member of $\overline{\textbf{10}}$, which was a suitable candidate for $N^{\ast}$ [15]. The state $N(1710)$, though listed in the PDG2002 baryon summary table [16] as a three-star resonance, was not seen in the PWA of the elastic scattering data $\pi N$ by the SAID group [17].
@@
Figure 3: Argand plots for the $P$ partial-wave amplitudes for the $\pi N$ case from threshold ($W=1080~{}\mathrm{MeV}$) to $W=2500~{}\mathrm{MeV}$ [19]. The crosses indicate $50~{}\mathrm{MeV}$ steps in $W$. The filled circles correspond to the BW $W_{R}$.
The $\pi$N PWA by the GW SAID group has shown that above the Roper resonance [18], $W>1500~{}\mathrm{MeV}$, $\sigma^{tot}\cong 2\sigma^{el}\cong 2\sigma^{inel}$ ($\sigma^{tot}=\sigma^{el}+\sigma^{inel}$). It means that it is almost pure diffraction, inelasticity $\eta\to\infty$ for $P_{11}$ and $S\cong 0$ with $A\cong i/2$. The result is that phase $\delta$ is poorly defined and the $P_{11}$ amplitude is spinning around $ReA=0$ [19] (Fig. 3).
Of course, the nonobservation of a broad $N(1710)$ state in $\pi N$ elastic analyses (no pole position, no BW, no speed plot) could be due to a very small $\pi N$ branching ratio. The standard procedure used in PWA may also miss narrow resonances with $\Gamma<20~{}\mathrm{MeV}$ by construction [17]. Therefore, the true non-strange unitary partner of $\Theta^{+}$ (if it is different from $N(1710)$ and sufficiently narrow) could have eluded detection.
Arndt and co-workers were considering the $\pi N$ partial wave $P_{11}$, as this is the amplitude that is associated with resonances having $J^{P}=1/2^{+}$ [20]. The character of $\chi^{2}$ changes, $\Delta\chi^{2}$, after inserting a narrow resonance into the partial amplitude with a range of masses, widths, and branching fractions is illustrated in Fig. 4. Negative values of $\Delta\chi^{2}$ emerge more readily near $M_{R}=1680~{}\mathrm{MeV}$ and $1730~{}\mathrm{MeV}$. We see that $\Delta\chi^{2}$ becomes negative only for $\Gamma_{el}=(\Gamma_{el}/\Gamma_{tot})\cdot\Gamma_{tot}$ within the bounds $\Gamma_{el}\leq 0.5~{}(0.3)~{}\mathrm{MeV}$ for $M_{R}=1680~{}(1730)~{}\mathrm{MeV}$ (Fig. 4). The available data cannot reliably discriminate values of $\Gamma_{el}$ below these bounds.
@@
(Fig. 5 (left)) [21]. BTW, in the case of a “neutron” target, the final-state interaction (FSI) is a problem. The effect is absent in the case of $\gamma p\to\eta p$ (Fig. 5 (right)) and agrees with the prediction of the Chiral Quark Solution approach [22]. In exact SU(3)F, the transition magnetic moment $\mu(p^{\ast}\to p)$ should vanish, since the U-spins are 3/2 for $p^{\ast}$, 1/2 for $p$, and 0 for the photon. With violation of SU(3)F, this transition moment does not vanish but is still much smaller than $\mu(n^{\ast}\to n)$ [23].
@@
The GW SAID group performed a coupled channel analysis of the $\pi N$ system that included elastic scattering of $\pi^{\pm}p$ and $\pi^{-}p\to\pi^{0}n$ with $\pi^{-}p\to\eta n$ [17]. The attractive factor for the reaction $\pi^{-}p\to\eta n$ is that it plays a role in isospin filtering. Unfortunately, the $\pi^{-}p\to\eta n$ data above $800~{}\mathrm{MeV}$ ($W=1630~{}\mathrm{MeV}$) are not reliable for PWA (see, for instance, Fig. 6 (left)). Most of the NIMROD data do not satisfy requirements (systematics is 10% or more, momentum uncertainties are up to $100~{}\mathrm{MeV/c}$, and so on). That is the main reason why the SAID coupled channel analysis had no opportunity to look for $N(1680)$ using reaction data $\pi^{-}p\to\eta n$ (Fig. 6 (right)).
Figure 6: Left: Excitation function for differential cross sections for the reaction $\pi^{-}p\to\eta n$. Data came from BNL, RHEL, and Saclay. Solid blue line from SAID PWA (WI08 solution) [26]. Right: Total cross sections for the reaction $\pi^{-}p\to\eta n$. Solid red line from SAID PWA [26]. All data are available at Ref. [25].
EPECUR Collaboration at ITEP performed high quality measurements of the differential cross sections for elastic $\pi^{\pm}p$ over an energy range of $p=800-1300~{}\mathrm{MeV/c}$ ($W=1.55-1.83~{}\mathrm{GeV}$) and for center-of-mass angles from 40 to $120~{}\mathrm{deg}$ [27]. In total, approximately 10,000 new data points have been obtained. These data have been produced with a momentum resolution of $\sim 1~{}\mathrm{MeV/c}$ and with $\sim 1$% statistical uncertainties.
Figure 7: Comparison of the measured $\pi$N data [25] and calculated $\pi^{-}p\to\eta n$ total cross sections. The solid red curve presents the calculations from the work [28]. The black dashed curve indicates the contribution of the $S$ wave.